If tan alfa =1,tan beta=2,tan gama=3,then alfa+beta+gama=

Grade:11

1 Answers

Arun
6 years ago
tan A = 1, tan B = 2, tan C = 3    (let alpha =A, beta = B, gama = C)
we know tan (A+B+C) = (tan A + tan B+ tan C – tanA *tanB *tanC)/ (1 – tanA *tanB – tanB *tanC – tanC *tanA)
now put the value,
tan(A+B+C) = (1 +2 +3 – 6)/(1 – 2 – 6 – 3)
tan(A+B+C) = 0/( – 10)
tan(A+B+C) = 0
hence
A+B+C = 0